OpenStudy (anonymous):

convert 21/40 to a percent

OpenStudy (anonymous):

take 21 divide it by 40 then multiply it by 100

OpenStudy (anonymous):

divide an move the decimal over twice \[\frac{21}{40+=.525\]

OpenStudy (anonymous):

i meant \[\frac{21}{40}=.525\]

myininaya (myininaya):

oh satellite your latex failed lol

OpenStudy (anonymous):

latex not working?

OpenStudy (anonymous):

lol

myininaya (myininaya):

the second one worked not the first one

OpenStudy (anonymous):

ok fine 21/40=.525=52.5%

OpenStudy (anonymous):

site is messing with me

OpenStudy (anonymous):

52.5*100?

OpenStudy (anonymous):

wat site is that anyways?

OpenStudy (anonymous):

no i multiplied by 100 by moving the decimal place twice to the right

OpenStudy (anonymous):

21/40=.525

myininaya (myininaya):

52.5%

OpenStudy (anonymous):

o i see

OpenStudy (anonymous):

21/40 = .525 .525 x 100= 52.5%

OpenStudy (anonymous):

then move decimal two spaces right and get 52.5%